OpenStudy (anonymous):

Find the Taylor series of 1/(1-x)^2 about a=0

OpenStudy (anonymous):

1+2 x+3 x^2+4 x^3+5 x^4+6 x^5+O(x^6)

OpenStudy (anonymous):

saubhik has it without finding derivatives, evaluating etc

OpenStudy (anonymous):

just by noting that \[\frac{1}{(1-x)^2}\] is the derivative of \[\frac{1}{1-x\]

OpenStudy (anonymous):

why dont you get alternating signs? when you differentiate 1/(1-x)^2 you get - sign

OpenStudy (anonymous):

oop i meant \[\frac{1}{1-x}\]

OpenStudy (anonymous):

yeah I get that part

OpenStudy (anonymous):

the well known formula for summing a geometric sequence \[\frac{1}{1-x}=1+x+x^2+x^3+...\]

OpenStudy (anonymous):

differentiate term by term to get the answer

OpenStudy (anonymous):

oh hell, I see... there are 2 minus signs

OpenStudy (anonymous):

oh in the chain rule? yes

OpenStudy (anonymous):

tripped me up the first time i saw it.

OpenStudy (anonymous):

:-) Now I get it, I just could not see why it is not 1-2 x+3 x^2-4 x^3+ and so on. but I see it now